A gym charges a fixed registration fee and an $8 fee for each class taken. It costs a total of $115 if one attends 5 classes.
A)write the equation in point-slope form to find the total cost y for taking among of classes
B)what is the fixed registration fee?
C)what is the total cost for a member that takes 20 classes at the gym?

Answer by TimothyLamb(4379)

b = fixed fee
---
answer A:
y = 8x + b
---
115 = 8*5 + b
b = 115 - 8*5
b = 75
---
answer B:
b = fixed fee = $75
---
answer C:
y = 8x + 75
y = 8*20 + 75
y = $235
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php